OpenStudy (anonymous):

(6t-4)=3t+1)+3 solve and check solution for t.

OpenStudy (anonymous):

6t - 4 = 3t + 4 3t = 8 t =8/3

OpenStudy (anonymous):

so what does t=

OpenStudy (anonymous):

t =8/3